A rectangular room is 1.8 times as long as it is wide, and its perimeter is 27 meters. Find the dimension of the room.

Answer by fractalier(6550)

Choosing L for length and W for width, the first statement reads
L = 1.8W
Then, for perimeter, we have
P = 2L + 2W = 27
We substitute and get
2(1.8W) + 2W = 27
3.6W + 2W = 27
5.6W = 27
W = 27/5.6 = 270/56 = 135/28
L = 1.8W = (9/5)W = (9/5)(135/28) = 243/28